# Nonlinearly charging the conformally dressed black holes preserving duality and conformal invariance

### v1

@@ Acknowledgements.
The authors would like to thank the NordGrav Summer Workshop recently held at the Instituto de Ciencias Exactas y Naturales, Universidad Arturo Prat, partially supported by the Agencia Nacional de Investigación y Desarrollo (ANID) under FONDECYT grant 11200742. D.F.-A. would like to thank the Departamento de Física of CINVESTAV-IPN for their kind hospitality and stimulating environment. We thank J. Barrientos, A. Cisterna, C. Corral, J. Oliva, and P.A. Sánchez for interesting discussions and suggestions. E.A.-B. has been partially funded by Conahcyt grant A1-S-11548. D.F.-A. is supported by ANID under grant FONDECYT No. 3220083. M.H. has been partially funded by FONDECYT grant 1210889.
@@ Appendix A General duality-invariant Lagrangians are necessarily implicit
which are warranted by (59b). Since this solution involves the single-argument arbitrary function $J$, it is in fact the general solution. Such implicit dependence generically characterizes all duality-invariant nonlinear electrodynamics in the Lagrangian formulation, see the recent results of Ref. [37] for a similar representation in different variables that we revise in a moment. It is important to emphasize that, as conditions (60) clearly indicate, the implicit dependence (59b) is strictly needed to have solutions only when the function $\Gamma$ is nontrivial. In the degenerate case where such function is a constant, (59b) is unnecessary and solution (59) becomes the standard sum separable solution. This degenerate sum-separable case is precisely the one corresponding to ModMax
@@
where the last relation implicitly determines the function $\Gamma(\mathscr{F},\mathscr{G})$ for each theory defined by the single argument function $J$. Being the only exception ModMax (61), as has already been explained.

### v2

@@ Acknowledgements.
The authors would like to thank the NordGrav Summer Workshop recently held at the Instituto de Ciencias Exactas y Naturales, Universidad Arturo Prat, partially supported by the Agencia Nacional de Investigación y Desarrollo (ANID) under FONDECYT Grant No. 11200742. D.F.-A. would like to thank the Departamento de Física of CINVESTAV-IPN for their kind hospitality and stimulating environment. We thank M. Astorino, L. Avilés, J. Barrientos, A. Cisterna, C. Corral, P. Fernandes, J. Oliva, and P.A. Sánchez for interesting discussions and suggestions. E.A.-B. has been partially funded by Conahcyt Grant No. A1-S-11548. D.F.-A. is supported by ANID under FONDECYT Grant No. 3220083. M.H. has been partially funded by FONDECYT Grant No. 1210889.
@@ Appendix A General duality-invariant Lagrangians are necessarily implicit
which are warranted by (59b). Since this solution involves the single-argument arbitrary function $J$, it is in fact the general solution. Such implicit dependence generically characterizes all duality-invariant nonlinear electrodynamics in the Lagrangian formulation; see the recent results of Ref. [38] for a similar representation in different variables that we revise in a moment. It is important to emphasize that, as conditions (60) clearly indicate, the implicit dependence (59b) is strictly needed to have solutions only when the function $\Gamma$ is nontrivial. In the degenerate case where such function is a constant, (59b) is unnecessary and solution (59) becomes the standard sum separable solution. This degenerate sum-separable case is precisely the one corresponding to ModMax
@@
where the last relation implicitly determines the function $\Gamma(\mathscr{F},\mathscr{G})$ for each theory defined by the single argument function $J$. The only exception is the ModMax theory (61), as has already been explained.